OpenStudy (anonymous):

hey there i need some here how to solve this Zxx-2Zxy+Zyy=(e^2x)cosy it's PDE i need steps to solve it especially the homogeneous part thanx

OpenStudy (anonymous):

\[_{Zxx} - _{2Zxy}+_{Zyy}=\exp \left( 2x \right) \cos y\]

OpenStudy (experimentx):

just a guess looks like x^2 - 2xy + y^2 --- for homogeneous part

OpenStudy (experimentx):

sorry ... i would like to correct that as x^2 - 4xy + y^2

OpenStudy (anonymous):

will it's a second order PDE i have the final solution it's : \[Z _{h}= f(x+y) +xf(x+y)\] but how to get there, that's what i need to know thanx for trying anyway ;)

OpenStudy (experimentx):

I've very little experience with PDE what is f(x+y)=? also the above seems to fit in PDE homogeneous part

OpenStudy (anonymous):

well tht's a part of what i didn't understand, i know f(x,y) , f is a function of x and y but this i havnt seen before

OpenStudy (anonymous):

:(

OpenStudy (experimentx):

stay tuned ... it's a good change for me to review ... i'll try to reply in few hrs

OpenStudy (anonymous):

what if the answer was really in f(x,y) instead of f(x+y) would it be possible to find the solution jst with a probability of a typing mistake in the note book i have lool ^_^

OpenStudy (anonymous):

any progress?